### v1

# Controllable Human-centric Keyframe Interpolation with Generative Prior

Abstract
Existing interpolation methods use pre-trained video diffusion priors to generate intermediate frames between sparsely sampled keyframes. In the absence of 3D geometric guidance, these methods struggle to produce plausible results for complex, articulated human motions and offer limited control over the synthesized dynamics. In this paper, we introduce PoseFuse3D Keyframe Interpolator (PoseFuse3D-KI), a novel framework that integrates 3D human guidance signals into the diffusion process for Controllable Human-centric Keyframe Interpolation (CHKI). To provide rich spatial and structural cues for interpolation, our PoseFuse3D, a 3D-informed control model, features a novel SMPL-X encoder that transforms 3D geometry and shape into the 2D latent conditioning space, alongside a fusion network that integrates these 3D cues with 2D pose embeddings. For evaluation, we build CHKI-Video, a new dataset annotated with both 2D poses and 3D SMPL-X parameters. We show that PoseFuse3D-KI consistently outperforms state-of-the-art baselines on CHKI-Video, achieving a 9% improvement in PSNR and a 38% reduction in LPIPS. Comprehensive ablations demonstrate that our PoseFuse3D model improves interpolation fidelity.
1 Introduction
Frame interpolation aims to generate new frames between two consecutive video frames to improve temporal smoothness. Traditional interpolation methods [11, 43, 16] assume small, simple motion over short time spans. These methods are challenged when the input frames are widely separated, known as keyframe interpolation or generative inbetweening [52, 44, 39, 10], where the motion between them becomes complex and ambiguous. This challenge is magnified in human-centric videos, where articulated body movements encompass diverse poses and shapes. With human subjects being prevalent in today’s video content, there is a growing need for interpolation methods to handle large temporal gaps and intricate human motion while offering plausible results.
Human-centric keyframe interpolation remains challenging for current methods. Recent approaches [10, 44, 39] leverage generative priors from image-to-video (I2V) models to bridge the temporal gap. These methods condition the interpolation solely on the input keyframes without intermediate guidance. Consequently, they often struggle to resolve motion ambiguities and accurately capture the complex articulated dynamics of human motion. For instance, when keyframes involve large occlusions or non-rigid joint movements, these methods often produce implausible or distorted interpolations (Figure 1(a)) due to insufficient intermediate guidance. FCVG [52] has explored interpolation keyframes with 2D skeletons as control signals for human subjects. However, 2D lines cannot convey full body shape and geometry, leading to unrealistic results (see Figure 1(b)). These methods lack fine-grained control over the interpolation process, limiting their ability to produce flexible, high-fidelity human-centric interpolations.
In this study, we investigate the integration of 3D human conditions into the human-centric keyframe interpolation pipeline. Drawing inspiration from recent advances in human animation [51, 50], we propose to integrate 2D human poses [45] with 3D SMPL-X models [28] as intermediate control signals. These signals provide precise guidance for complex articulated motions: 2D poses offer concise representations of human joint poses, while 3D models capture rich spatial geometry. However, effectively processing these control signals poses challenges. First, common practice renders 3D human models into 2D proxies (e.g., colored surface, normals, depth maps) before encoding, leading to substantial loss of spatial information in occluded regions. Therefore, we need to develop a dedicated encoder that preserves occluded 3D details when converting models into control signals. Second, fusing signals with different information content and granularity is nontrivial. This necessitates designing appropriate neural architectures that can accurately extract 3D cues and harmonize them with 2D poses into a unified, informative control input.
Figure 1: Keyframe Interpolation with Different Strategies. (a) Interpolation using I2V models without intermediate guidance often yields implausible or distorted frames, especially under large motion or occlusion. (b) Skeleton-guided interpolation offers structural cues but lacks geometric detail, resulting in unrealistic body shape and appearance. (c) Our PoseFuse3D-KI employs dense human-centric guidance, enabling temporally coherent and visually plausible interpolations.
To this end, we introduce PoseFuse3D Keyframe Interpolator, termed as PoseFuse3D-KI, a novel pose-control framework for controllable human-centric keyframe interpolation (Figure 1(c)). Our framework is unique in its 3D-informed control model PoseFuse3D, which comprises three jointly trained modules. The first derives control features from visualized conditions; the second encodes and aggregates 3D body geometry into 2D image conditioning features; and the third combines outputs from the first two modules into a unified control signal for interpolation. In contrast to Champ [51], which fuses 3D features from rendered visualizations, our encoder processes features directly in 3D and integrates projected features through feature aggregation.
To evaluate the proposed PoseFuse3D-KI, we have also built a high-quality video dataset for Controllable Human-centric Keyframe Interpolation (CHKI). Most existing interpolation datasets, such as SportsSlomo [5], target small temporal gaps, lack annotations for 2D poses or 3D human models, and offer limited human-centric motion diversity. Therefore, we introduce CHKI-Video, a new dataset for a systematic evaluation of CHKI algorithms. CHKI-Video comprises 2,614 high-quality video clips of over 180K frames sourced from SportsSlomo [5] and Pexels [^1] https://www.pexels.com/videos/ website that hosts high-quality stock videos. Each frame is carefully annotated with bounding boxes, segmentation masks, 2D human poses, and SMPL-X parameters, using state-of-the-art tools [3, 18, 23, 32, 40] supplemented by manual verification. From this collection, we derive a benchmark specifically for the CHKI task. We hope this benchmark will help improve controlled keyframe interpolation techniques with its high-quality videos and diverse examples.
Our contributions are threefold: (i) We present PoseFuse3D-KI, an effective interpolation framework for human-centric keyframe interpolation, characterized by a novel pose-control model, PoseFuse3D. It effectively extracts control signals from 3D SMPL-X and fuses 2D signals, allowing precise and informative control. (ii) For evaluation, we construct CHKI-Video, a benchmark dataset with comprehensive human-centric annotations, which are absent in existing interpolation benchmarks. (iii) Through extensive experiments, we demonstrate that PoseFuse3D-KI delivers state-of-the-art performance on our CHKI-Video benchmark with an improvement of 1.85 dB in PSNR and a reduction of 0.0796 in LPIPS.
2 Related Work
Frame Interpolation. Traditional frame interpolation methods are primarily designed for temporally adjacent frames and rely on either direct synthesis using convolutional networks [7, 17], or motion representations such as dynamic kernels [26, 27, 30, 6, 8, 20] and optical flows [15, 21, 19, 35, 13, 1, 11, 16, 41, 46]. Recent advances [10, 44, 39] extend this task to more challenging keyframe interpolation scenarios by leveraging the generative priors of image-to-video diffusion models [2, 37, 33, 12]. These methods combine temporal forward and reverse denoising predictions in a unified process to enable interpolation. However, such approaches struggle when faced with complex articulated motions or ambiguous transitions from human keyframes. To alleviate motion ambiguity in interpolation, FCVG [52] introduces 2D matched lines as control signals. However, this control signal lacks the 3D geometric context required for plausible human-centric interpolation, resulting in unrealistic body shape and appearance.
Pose-Guided Human Animation. Recent advances in pose-guided human animation [50, 51, 14, 31, 47, 42, 38] harness the power of diffusion models and have achieved remarkable success in generating videos from a single reference image. These methods offer flexible and precise control by incorporating enriched conditioning signals and increasingly sophisticated control mechanisms. 2D human poses, such as OpenPose [4] and DWPose [45], are widely used in existing works [51, 50, 31, 47], but they are limited in capturing fine-grained geometry and motion dynamics. To address this, recent works [51, 50] integrate 3D human parametric models [25, 28], which offer realistic body representation through blend shapes and skinning, resulting in more accurate and expressive human animations. Specifically, Champ [51] renders 3D human models into 2D proxies (e.g., normals, depth, and semantic maps) and combines them with 2D pose visualizations as control input. It operates directly on these visualizations and unifies them using a simple summation operation. To incorporate control signals into diffusion models, many approaches [38, 42, 47, 50] adapt ControlNet [47] for their customized control networks. Some methods [51, 14] introduce task-specific pose guiders but often require retraining the majority of denoiser parameters. ControlNeXt [31] improves control efficiency by encoding conditions with a lightweight convolutional network and injecting them via cross-normalization after the first denoising block, tuning only a minimal subset of parameters. This efficient mechanism enables robust conditioning over large-scale pre-trained video generators [37, 2].
In this paper, we demonstrate the advantage of combining 2D poses (DWPose [45]) with 3D human models (SMPL-X [28]) through comprehensive experiments. We present a novel pose control model that extracts unified, 3D-informed control features to provide precise guidance. Instead of relying on rendered normals and depth maps, the model directly extracts explicit 3D information from human models in 3D space, preserving richer control signals. It adopts a ControlNeXt-inspired strategy to control video diffusion models for keyframe interpolation.
3 Method
Given a human-centric keyframe-pair $I_{0},I_{N}\in\mathbb{R}^{H\times W\times 3}$ with timesteps $\{0,N\}$, Controllable Human-centric Keyframe Interpolation (CHKI) is formulated as:
$\{\hat{I}_{i}\}^{N-1}_{i=1}=\mathcal{G}(I_{0},I_{N},\{\textbf{C}\}^{N}_{i=0})$ (1)
where $\mathcal{G}$ denotes an interpolation model guided by control signals $\{\textbf{C}_{i}\}$. In this work, we aim to address CHKI by proposing an effective controllable interpolation framework, PoseFuse3D-KI. This framework integrates 3D-aware human-centric signals into a pre-trained Video Diffusion Model (VDM) through our 3D-informed control model PoseFuse3D, as illustrated in Figure 2(a).
3.1 PoseFuse3D
PoseFuse3D is a 3D-informed control model that provides 3D human structure and geometry guidance for plausible human interpolation. This 3D-informed guidance is injected into the base diffusion model after the first denoising block via cross-normalization [31]. Internally, PoseFuse3D comprises three jointly trained components: a visual encoding module that derives control features from 3D SMPL-X [28] renderings and 2D DWPose [45] visualizations; a SMPL-X encoder that directly embeds 3D humans and aggregates them into image conditioning maps; and a fusion module that integrates encoding streams into a unified control tensor to guide interpolation in the base VDM.
Figure 2: Model Architecture. Our PoseFuse3D-KI framework, as shown in (a), comprises a video diffusion model (VDM) and a novel control model, PoseFuse3D. The PoseFuse3D model extracts rich features from both 3D and 2D control signals and fuses them into a unified representation to guide the VDM. The key component of PoseFuse3D is the SMPL-X encoder as illustrated in (b), which provides explicit 3D signal features. Specifically, the SMPL-X encoder first extracts 3D information from the SMPL-X model with 2D correspondences via projection. The 3D and 2D information is then encoded in parallel. With features of 2D correspondences, 3D information is aggregated onto the 2D image plane using attention mechanisms. The aggregated features are subsequently processed to produce the final feature $S^{3D}$.
Visual Encoding. This visual encoding module extracts conditional features from visualized control images, maintaining natural pixel-level alignment with the controlled video latents. To enhance the control signals, we incorporate visualizations from both DWPose and SMPL-X. While SMPL-X renderings provide rich human surface details, their keypoint information is indirect, mixed with other vertices and mesh faces. Therefore, we add DWPose visualizations to emphasize the skeletal keypoint layout for robust pose understanding. This combination was also demonstrated to be effective in prior work [51]. Specifically, our visual encoding module employs two parallel convolutional encoders to capture comprehensive pose information. One encoder processes DWPose visualizations to capture compact pose information, while the other handles SMPL-X renderings that retain 3D cues such as occlusion boundaries and projected shapes. Notably, to enrich semantic detail, we use SMPL-X Colored Surface [50] during rendering, which assigns a unique color to each vertex. The resulting feature maps, $E^{2D}$ and $E^{3D}$, are then passed to the fusion module for unified conditioning.
SMPL-X Encoder. Although 3D model renderings offer aligned image maps for conditioning in VDMs, the rendering operation discards parts of 3D information, particularly in occluded regions. This results in implausible interpolation for keyframes of large human motion. For example, in the first case of Figure 3, all variants relying solely on rendered controls fail to interpolate the occluded arms with correct spatial position. To enhance controllability with direct 3D information, we introduce a SMPL-X encoder (Figure 2(b)) that processes the SMPL-X model in 3D space and transforms it into image conditioning feature $S^{3D}$. Specifically, a SMPL-X model [28] is parameterized with $(\theta,\beta,\psi)$, corresponding to pose, shape, and expression. We obtain structural information by forwarding these parameters into the SMPL-X model to generate vertex and joint coordinates $P^{3D},J^{3D}$ in 3D space and obtain their corresponding 2D coordinates $P^{2D},J^{2D}$ through projection:
$P^{3D},J^{3D}=\mathrm{SMPLX}(\theta,\beta,\psi),$ (2)
$P^{2D},J^{2D}=\mathrm{Projection}(P^{3D},J^{3D}),$ (3)
where the joint coordinates $J^{3D},J^{2D}$ correspond to the pose parameter $\theta$ that indicates joint rotations. Notably, the projection step establishes a correspondence between the 3D space and the 2D image plane, making it possible to retain 3D spatial structure while producing image conditioning maps. Next, the raw 3D coordinates are processed using an MLP to produce point-wise vertex features $V^{P}$ and joint features. These joint features are refined through a temporal residual block that fuses them with pose information into expressive joint-level representations $V^{J}$. To aggregate these 3D features into 2D image control maps for joints and vertices, we employ separate attention mechanisms:
$O^{J}=\mathrm{JointAttn}(Q,K^{J},V^{J}),$ (4)
$O^{P}=\mathrm{VertexAttn}(Q,K^{P},V^{P}),$ (5)
where $Q\in\mathbb{R}^{B\times HW\times d}$ is the flattened $d$-dimension feature of a standard 2D meshgrid extracted by a convolutional encoder. Finally, the outputs $O^{J}$ and $O^{P}$ are concatenated and passed through a downsampling block to produce the final SMPL-X control representation $S^{3D}$, which serves as an informative and compact image embedding of the underlying 3D human structure.
Condition Fusion. The condition fusion module combines control features from both 2D and 3D signals into a unified representation to guide keyframe interpolation. For robust feature representation, we introduce a coarse-to-fine fusion strategy that progressively integrates rich geometric information from the 3D features into the compact 2D pose features. Specifically, we adopt two attention-based fusion blocks to perform this integration, where each block contains three attention layers for progressive refinement. The first layer is a self-attention module that processes the 3D features by operating on the sum of rendering encoding $E^{3D}$ and SMPL-X features $S^{3D}$. The second layer performs cross-attention, aligning the 3D features with the 2D encoding $E^{2D}$ through a spatially localized interaction scheme. Notably, we adapt the shifted window-partition strategy [24, 46] to restrict attention computation to adjacent regions, enhancing local alignment. The third layer applies temporal self-attention to capture temporal dynamic correlations in the fused representation. We use the second fusion block output as the final control signal, which is injected into the base interpolation engine to provide fine-grained, structure-aware guidance during synthesis.
3.2 VDMs as Base Interpolation Engine
To supply generative priors for human-centric keyframe interpolation, we adapt pre-trained Video Diffusion Models (VDMs) under the latent diffusion framework [33]. VDMs perform the diffusion process in a VAE-encoded latent space, conditioning video synthesis on input frames. Our main experiments investigate Wan2.1 [37] as the base model for keyframe interpolation.
Wan2.1 consists of a scaled-up DiT-based denoiser [29] and a causal 3D-VAE that performs spatiotemporal compression. It employs the flow matching strategy [22, 9] for the diffusion process. It formulates the forward diffusion process as a linear interpolation between the clean video latent z and the noise $\bm{\epsilon}$, which adds the noise by: $\textbf{z}_{t}=(1-t)\textbf{z}+t\bm{\epsilon}$. In the backward process, the denoiser $f_{\theta}$ iteratively refines $\mathbf{z}_{t}$ conditioned on the first frame $I_{0}$. The training objective is formulated as:
$\mathcal{L}=\mathbb{E}_{\textbf{z},t,I_{0},\bm{\epsilon}\sim\mathcal{N}(0,I)}[||f_{\theta}(\textbf{z}_{t},I_{0},t)-\textbf{y}||_{2}^{2}],$ (6)
where the target objective y is $\frac{d\textbf{z}_{t}}{dt}=\bm{\epsilon}-\textbf{z}$. Since the latent space is unevenly compressed in time, noise fusion strategies [10, 52] that combine temporally forward and reverse denoising paths are ineffective. Instead, we adapt Wan2.1 for keyframe interpolation in one unified denoising process. Specifically, we condition Wan2.1 on both endpoint frames $I_{0},I_{N}$, and apply LoRA tuning to its input patch embeddings. Additional details are provided in the supplementary material.
4 The CHKI-Video Dataset
Existing interpolation datasets [43, 7, 5, 34, 36], which are designed for interpolating temporally adjacent frames, are not suitable for the CHKI task. To address this gap, we built the CHKI-Video dataset in three consecutive stages. More details can be found in the supplementary material.
Stage 1: Dataset Collection. To create a dataset with challenging and diverse human motion, we curate video clips from SportSlomo [5] and Pexels [^2] https://www.pexels.com/videos/. While SportsSlomo provides challenging human-centric videos, its exclusive focus on sports limits the diversity of activities. To enhance diversity, we compile a list of keywords, spanning everyday activities to high-intensity actions. We use these keywords to retrieve videos from Pexels, complementing the sports videos. The curated videos are then downsampled to eliminate frame redundancy unnecessary for keyframe interpolation.
Stage 2: Pre-annotation Processing. We first perform general filtering for the low-quality videos according to brightness changes and assessed scores [40]. Then, we use Grounding-DINO [23] and SAM2 [32] to detect, segment, and track human instances in each video. We discard any video of more than three people or is shorter than 20 consecutive frames to guarantee a sufficient temporal span for keyframe interpolation. After automated processing, we manually review and filter detections in complex sports scenarios.
Stage 3: Human-centric Annotation. Building on the accurate human detections obtained in Stage 2, we annotate each clip for precise human-centric information. First, we employ Sapiens [18] to extract 2D human keypoints and perform whole-body detection to filter out clips with incomplete human figures. This ensures our dataset remains strictly human-centric. Finally, we apply SMPLer-[3], leveraging its high re-projection accuracy to fit detailed SMPL-X models for human images and produce reliable 3D body parameters for each frame.
As a result, CHKI-Video comprises 2,614 video clips of over 180K frames, carefully annotated with human bounding boxes, masks, 2D keypoints and 3D parametric models. To prepare the train and test split, we follow the original division for SportsSlomo [5] videos and distribute the Pexels videos according to their keyword frequencies to maintain balanced coverage of all motion categories.
5 Experiments
We present quantitative and qualitative results in Sec. 5.1 to validate the effectiveness of our 3D control strategy in PoseFuse3D. We compare our interpolation performance against state-of-the-art methods in Sec. 5.2. Finally, Sec. 5.3 provides a detailed ablation study to justify our model design.
Implementation Details. We fine-tune PoseFuse3D-KI on the CHKI-Video training split for 70k iterations. Specifically, we fine-tune our 3D-informed control model PoseFuse3D, and employ LoRA on the input patch embeddings, as well as the value and output projections of the VDM’s attention modules. During training, we randomly sample 25 consecutive frames from video clips and process them to a resolution of $512\times 320$. To maximize human-centric content, we crop each clip around its largest annotated human bounding box. Given the ratio from the target input size, we first crop the maximum scale of the image with the largest box as the center. The cropped image is then resized to match the target resolution. More implementation details are provided in the supplementary material.
Evaluation Protocols. We evaluate methods on CHKI-Video’s test set using the ground-truth annotations to drive controllable interpolation. With motion ambiguity limited by the ground truth controls, we adopt the standard interpolation metrics of PSNR and LPIPS computed on the whole images for evaluation. To quantify performance specifically on human regions, we further leverage the annotated human boxes and masks, yielding PSNRbbox, LPIPSbbox, PSNRmask, and LPIPSmask metrics. Notably, we apply binary dilation to expand and smooth the masks before computing the metrics, preventing potential artifacts along mask boundaries.
5.1 3D Control Strategy in PoseFuse3D
Table 1: Comparisons of 3D Control Strategies. This table presents quantitative results of different 3D control strategies. The PSNR and LPIPS metrics are calculated for the whole image, as well as for the human-centric parts via annotated boxes or masks. Best results are highlighted with boldface.
Method Backbone 3D Control Strategy Evaluation Metrics
[…]
FCVG SVD N.A. 20.42 13.28 12.14 0.2100 0.0899 0.0606
PoseFuse3D SVD VE-SVD 20.96 13.79 12.70 0.1975 0.0835 0.0553
PoseFuse3D Wan2.1-I2V VE 21.91 14.36 13.10 0.1400 0.0682 0.0484
PoseFuse3D Wan2.1-I2V VE+DN 22.07 14.35 13.03 0.1363 0.0667 0.0473
PoseFuse3D Wan2.1-I2V VE+SE (Ours) 22.14 14.53 13.24 0.1330 0.0653 0.0464
Figure 3: Qualitative Results of Different 3D Control Strategies. We use red circles to highlight regions where the 3D controls and our strategy significantly improve the interpolation quality.
Setup. We evaluate the effectiveness of the 3D-informed control design in PoseFuse3D via comparisons among 3D control strategies, including ‘VE’, ‘VE+DN’ and ‘VE+SE’. Specifically, VE refers to the visual encoding on the SMPL-X colored surface rendering, removing the SMPL-X encoder from the full design of our control model. VE+DN extends VE by incorporating depth and normal renderings, using two additional encoders identical to the one used in VE. VE+SE represents our proposed strategy in PoseFuse3D that directly encodes 3D information via the SMPL-X encoder and integrates it with VE. Experiments for these strategies use Wan2.1-I2V [37] as the interpolation backbone. For efficiency, they are trained for 40K iterations.
To assess the necessity of 3D information, we compare against FCVG [52], which is conditioned solely on 2D signals. For fair comparison, we create a variant of VE by replacing the backbone with SVD [2]. We adapt SVD for interpolation with the temporal forward and reverse denoising path fusion strategy used in FCVG. This variant is referred to as VE-SVD for ease of analysis.
Results. We provide quantitative comparisons of 3D control strategies in Table 1.In comparisons between FCVG and VE-SVD, we find that adding 3D control improves the interpolation performance. VE-SVD outperforms FCVG across metrics with a more than 0.50 dB increase on all the PSNRs, indicating the improvements on both whole-image and human-centric levels. Moreover, our study highlights the importance of explicit 3D information. VE+DN and VE+SE, which incorporate depth and normal maps or direct SMPL-X information, outperform the simpler strategy VE. VE+DN and VE+SE show clear improvements in perceptual quality as reflected by the LPIPS metrics. Notably, our VE+SE strategy, which directly encodes information in 3D space, delivers the best performance, achieving the lowest LPIPSbbox of 0.0653 and the highest PSNR of 22.14 dB.
Visualizations. Besides our quantitative analysis, we perform a qualitative comparison of the 3D control strategies, as shown in Figure 3. We find that incorporating 3D controls better preserves human shape during interpolation. Take the Tennis case in Figure 3 for example, methods with 3D control strategies interpolate the player’s body close to the ground truth, whereas FCVG exhibits noticeable distortion. Moreover, our VE+SE, which directly encodes 3D information from SMPL-X, proves effective in handling occluded human motion. In both the Skateboarding ($1^{st}$ row) and Jumping ($2^{nd}$ row) cases in Figure 3, we can observe that our VE+SE strategy produces plausible results for the occluded arms, demonstrating its advantage in complex scenarios.
5.2 Benchmark Results
Table 2: Comparisons with State-of-the-art Interpolation Methods.
Methods Metrics
[…]
GIMM-VFI [11] 20.29 11.59 10.16 0.1954 0.1187 0.0860 0.9146
GI [39] 15.81 8.27 7.26 0.3364 0.1672 0.1146 0.8954
Wan2.1-KI (Ours) 19.53 11.19 9.85 0.2081 0.1208 0.0868 0.9180
FCVG [52] 20.42 13.28 12.14 0.2100 0.0899 0.0606 0.9187
PoseFuse3D-KI (Ours) 22.27 14.72 13.47 0.1304 0.0636 0.0450 0.9189
Table 3: Ablation Study on Visual Encoding and Fusion Module.
Model Variants Evaluation Metrics
[…]
Conv-Enc (2D) N.A. 20.29 12.17 11.33 0.1990 0.1096 0.0836
Dual Conv-Enc Sum 20.50 12.43 11.59 0.1953 0.1059 0.0806
Dual Conv-Enc Non-TSA 20.30 12.19 11.35 0.1956 0.1053 0.0790
Dual Conv-Enc Non-WP 20.46 12.42 11.60 0.1938 0.1038 0.0775
Dual Conv-Enc Full 20.55 12.53 11.71 0.1927 0.1031 0.0773
Figure 4: Qualitative Comparisons with State-of-The-Art Methods.
Setup. We compare PoseFuse3D-KI against several advanced interpolation methods on our CHKI-Video dataset. The main comparison is with FCVG [52], which also enables intermediate control during interpolation. For broader coverage, we also include the keyframe interpolation method GI [39] and the traditional video frame interpolation method GIMM-VFI [11]. We also include Wan2.1-KI, an adaptation of Wan2.1 [37] for keyframe interpolation, following the strategy in Sec. 3.2. Since not all baselines support ground truth controls, we additionally compute the Human Anatomy (HA) [49] score to assess the quality of human synthesis and support fair comparison.
Results. We report quantitative results of PoseFuse3D-KI on the CHKI-Video benchmark in Table 2. Our method delivers state-of-the-art performance on human-centric keyframe interpolation. On whole-image metrics, it boosts PSNR by 1.85 dB and lowers LPIPS by 0.0796 in comparison with the state-of-the-art method FCVG [52]. Crucially, it also outperforms on human-centric metrics compared with other methods, achieving a PSNRbbox of 14.72 dB, an LPIPSmask of 0.045, and an HA score of 0.9189. This indicates that our method produces plausible, high-fidelity human interpolations that closely follow the ground-truth dynamics, demonstrating the effectiveness of our method.
Visualizations. For qualitative evaluation, we qualitatively compare PoseFuse3D-KI with other advanced methods, as illustrated in Figure 4. Our approach delivers more accurate human interpolations, faithfully following real-world motions and preserving body shape. For example, in the $2^{nd}$ ‘Fencing’ case and $4^{th}$ ‘Stunt Bike’ case, only PoseFuse3D-KI correctly interpolates leg and arm movements while maintaining consistent shapes. Moreover, our method naturally handles the occluded human motion well with 3D-informed control. We observe that our method interpolates correct spatial positions for the occluded legs ($1^{st}$ case) and arms ($3^{nd}$ case), achieving significant improvements over FCVG. Furthermore, although the control-free keyframe interpolation methods GI and Wan2.1-KI occasionally generate undistorted humans, they often generate implausible motion that violates real-world dynamics, as observed in $1^{st}$, $2^{nd}$, and $4^{th}$ cases.
5.3 Ablation Study
In this section, we ablate the visual encoding and fusion modules of PoseFuse3D in Table 3. For efficiency, we use SVD as the backbone and process video clips into 9 consecutive frames of $256\times 256$. We use 3× temporally downsampled videos from the CHKI-Video test set for evaluation.
3D Visual Encoding. Our visual encoding module includes two convolutional encoders for 2D and 3D control maps, respectively. We denote the variant including this entire module as ‘Dual Conv-Enc’, and the one using only the 2D encoder as ‘Conv-Enc (2D)’. Removing the 3D visual encoding leads to a performance drop of 0.26 dB on both PSNRbbox and PSNRmask, highlighting the importance of 3D visual encoding.
Fusion Module. In PoseFuse3D, condition features are fused through a carefully designed fusion module. To validate its effectiveness, we replace it with a simple summation operation [51], denoted as ‘Sum’ in Table 3. This change leads to a significant performance drop, particularly in perceptual quality, with an increase of 0.0033 in LPIPSmask. These results demonstrate the fusion module’s contribution to providing informative control for high-quality interpolation.
Window-Partition Strategy. PoseFuse3D employs a cross-attention layer with a shifted window-partition strategy to fuse features across neighboring windows. To validate this design, we remove the window partitioning, denoted as Non-WP. This results in notable drops of 0.11 dB in both PSNRbbox and PSNRmask, indicating that the window-partition strategy enhances controlled interpolation.
Temporal Attention in Fusion Module. To justify the efficacy of the temporal self-attention (TSA) layer in the fusion module, we conduct experiments excluding the TSA layer (Non-TSA). This removal causes increases of 0.0022 and 0.0017 in LPIPSbbox and LPIPSmask, demonstrating the crucial role of the temporal self-attention layers in the fusion module.
6 Conclusion
We propose PoseFuse3D-KI, a controllable human-centric keyframe interpolation framework powered by our novel 3D-informed control model, PoseFuse3D. PoseFuse3D embeds rich spatial geometry from 3D human signals together with 2D poses into a unified control feature, enabling the generation of more plausible and realistic intermediate frames. For evaluation, we construct a CHKI-Video dataset with comprehensive human-centric annotations. Extensive experiments on the benchmark demonstrate that PoseFuse3D-KI outperforms previous interpolation methods with a 9% improvement in PSNR and a 38% reduction in LPIPS.